# Beyond Film Subtitles: Is YouTube the Best Approximation of Spoken Vocabulary?

### v1

@@ 1 Introduction
Word frequency is crucial for psycholinguistic research, as well as for assistive or educational applications involving production or comprehension of words. Traditionally, written corpora have been used for estimates of word frequency, with Kučera and Francis (1967) frequency norms long dominating psycholinguistic research of English.
The size of available written language corpora has grown over time, while speech corpora are still costly to develop and comparably limited in extent. When user-generated text became available on a large scale, it was possible to approximate everyday language exposure by collecting English text from Usenet newsgroups (Burgess and Livesay, 1998), and later French film and TV subtitles (New et al., 2007). Subtitle-based norms for US English, SUBTLEX-US, (Brysbaert and New, 2009) were found more predictive of lexical decision times (LDT) than frequencies based on traditional written corpora or the Usenet-based corpus.
@@
As full corpus data cannot be published due to copyright, we release two basic language models based on the TUBELEX corpus for each language in addition to the frequency lists: a statistical language model (Heafield et al., 2013), which provides smoothed frequencies of word 1-grams to word 5-grams, and fastText word embeddings (Bojanowski et al., 2017) to enable modeling of semantic similarity or analogy, as well as representation of words in downstream application. FastText extends the Word2vec model (Mikolov et al., 2013). Preprocessing details and hyperparameters are provided in Appendix A, model sizes in Appendix B, and evaluation of the embeddings in Appendix C.
@@ 2.1 Subtitle Corpora
All of these corpora were built with the intent of approximating spoken language, and most of them were evaluated against psycholinguistic data. Other subtitle corpora were built for different purposes:
@@ 3.2 Cleaning and Duplicate Removal
We identify the language of each subtitle line using the compressed fastText language identification model444https://fasttext.cc/docs/en/language-identification.html (Joulin et al., 2016a, b). We discard files containing less than 95% of the target language. From the remaining files, we remove both lines that do not contain any valid characters for the target language (e.g. Latin alphabet for English), and lines that are identified as a different language. Lastly, we discard any files consisting of less than three lines of text. The resulting numbers of files are listed as Cleaned in Table 1.
@@ 3.4 Masking Personal Information
We also use special tokens to replace sequences of digits (after tokenization) and anonymize email addresses, web addresses including those without an explicit protocol (e.g. x.com/username), and apparent social network handles starting with @. Our approach to anonymizing personally identifying information (PII) is informed by the analysis by Subramani et al. (2023) and extends the previous approach of (Soldaini et al., 2024) by also masking web addresses and social network handles.
@@ 3.5 Tokenization and Frequency Lists
English, Indonesian, and Spanish segmented using Stanza (Qi et al., 2020) tokenize,mwt pipeline ; Japanese segmented using MeCab (Kudo et al., 2004), and the UniDic 2.1.2 dictionary (Den et al., 2007, distributed as unidic-lite777https://pypi.org/project/unidic-lite/); Chinese segmented using the jieba888https://github.com/fxsjy/jieba segmenter 0.42.1.
@@
| Evaluation Label | Chinese | English | Indonesian | Japanese | Spanish |
| Subtitles match speech in the target language | 65% | 91% | 84% | 84% | 89% |
| Subtitles match song in the target language | — | 2% | 1% | 1% | 2% |
| Audio description | — | 1% | 1% | 0% | 0% |
| No speech or song | — | 2% | 0% | 0% | 2% |
| Synthesized speech | 34% | 3% | 13% | 14% | 5% |
| Audio language differs | 1% | 1% | 1% | 1% | 2% |
| Subtitle language differs | — | — | — | — | — |
@@ 4 Human Evaluation
We performed human evaluation to verify how representative the corpus is of the target languages, and spoken language in particular. We took a sample of size 300 for each language, consisting of 100 videos with 3 random subtitle time stamps for each. The videos were selected using stratified sampling by category and duration999 We divided the videos in three similarly large duration classes: [0, 3 min), [3 min, 10 min), and [10 min, ). for each language. Each sampled timestamp was examined and labeled by a CEFR C2-level non-native speaker for English and by native speakers for the other languages.
The labels and evaluation results are shown in Table 2. Most importantly, among the 1,500 subtitle cues, we have not found a single one whose language would differ from the target language. We have, however, observed different dialects or varieties of each language, as well as apparent non-native speech, sometimes co-occurring in the same video. In the case of Chinese, 2 of the 100 videos were in Cantonese, with traditional Chinese subtitles, while the majority was in Mandarin Chinese. To better understand the composition of the Chinese subtitles, we also analyzed the script used in the whole corpus, and found that 66% videos use simplified Chinese, 33% videos use traditional Chinese, and 1% mix both.101010 We used the Hanzi Identifier package (https://github.com/tsroten/hanzidentifier). and considered subtitles mixed if they contained the non-majority script variant on at least 2 lines and at least 5% of lines.
@@ 5.1 Evaluated Corpora and Resources
GINI, a Twitter-based metric (Murayama et al., 2018), measuring words’ dispersion in frequency of use by different people; Wikipedia; wordfreq, a Python library (Speer, 2022) pooling frequency from multiple corpora. Wordfreq combines Wikipedia, Twitter and a subtitle corpus for each of the evaluated languages, as well as 4 more sources for English, Chinese, and Spanish, and 2 more for Japanese. The subtitle data used by wordfreq is OpenSubtitles2018, SUBTLEX-US and SUBTLEX-UK for English, SUBTLEX-CH for Chinese.
For each corpus, we provide technical details in Appendix D, and token and type counts in Appendix E.
@@ 5.2 Computing Frequency
where is the number of occurrences of the word , #tokens is the total number of tokens in the corpus, and #types is the number of types in the corpus. As a result, even words missing in the corpus are assigned a non-zero frequency.
For ACTIVE-ES and wordfreq, which do not provide token counts, we instead assign the corpus minimum frequency to missing words. We also directly use GINI values, analogously assigning the corpus maximum value to missing words, as high GINI values indicate high dispersion.
The words featured in our experiments may be tokenized as multiple tokens. We assign them the minimum of the frequencies of the individual tokens, and maximum of the GINI values.
In all experiments we use the logarithm of the above values, except for GINI, where it consistently worsened performance. For a GINI value , we use the additive inverse for the purpose of comparison with log-frequencies.
@@ 5.3 Lexical Decision Time
Other corpora either performed comparably, but only covered a single language (SubIMDB, EsPal) or underperformed noticeably in at least one language (SUBTLEX in Spanish, wordfreq in Chinese, Wikipedia in all languages, and GINI and ACTIV-ES in the one language they cover). Furthermore, as can be seen in Figure 1, TUBELEX and SUBTLEX perform remarkably well relative to their size.
@@ 5.4 Word Familiarity
We measure correlation (PCC) with mean word familiarity from five databases: Chinese familiarity ratings (Su et al., 2023), English MRC lexical database (Coltheart, 1981; Coltheart and Wilson, 1987), Indonesian lexical norms (Sianipar et al., 2016), Japanese word familiarity ratings (Asahara, 2019)111111 We use the published ratings for reception estimated using a Bayesian linear mixed model. , and Spanish lexical norms (Guasch et al., 2016). Evaluation on three alternative, smaller databases for English, Spanish, and Japanese can be found in Appendix F.
As shown in Table 4, in each language except Japanese, TUBELEX’s correlation is either the strongest one or not significantly weaker. Correlations without any significant difference from TUBELEX are achieved by SUBTLEX in Chinese, by all subtitle corpora and SubIMDB and wordfreq in English, by wordfreq in Indonesian, and by all subtitle corpora except EsPal, and by CREA-Spoken in Spanish. In Japanese, GINI achieves the strongest correlation, followed by TUBELEX and TUBELEX.
@@ 5.5 Lexical Complexity
As the dataset was used for evaluation of lexical complexity prediction in a shared task (Shardlow et al., 2024), we also compare predictions based on TUBELEX with top shared task participants. To do so, we fit a linear regression model using a single variable, log-frequency or GINI values to the shared task’s trial data (30 instances for each language), and clip the predicted values to the range [0, 1]. We compare our results with the shared task submissions that achieved the highest coefficient of determination (and also highest correlation) in individual languages (TMU-HIT, Enomoto et al., 2024; and GMU, Goswami et al., 2024), and on average across all shared task languages (Archaeology, Cristea and Nisioi, 2024).
The best results are achieved by TUBELEX, closely followed by TUBELEX, in English; TUBELEX, closely followed by TMU-HIT and TUBELEX, in Japanese; and by TMU-HIT in Spanish, where it outperformed others by a large margin. As we are using cited values achieved by task participants, we cannot evaluate statistical significance in this case.
@@
It should be noted that, in this case, we are evaluating the prediction of lexical complexity, not a mere correlation with it. If we looked only at correlation, thus ignoring misprediction of mean and variance, TMU-HIT’s predictions would be more strongly correlated than TUBELEX’s in the three languages, and those of the other two systems in English (complete results provided in Appendix H). This may indicate limitations of LLM prompting as a regression method.
@@ 6.1 Tokenization and Lemmatization
For Japanese the base form performs the best for familiarity, and default tokenization for non-native lexical complexity. Both always outperform the orthographically normalized lemma, which show the importance of the exact written form in Japanese.
@@ 6.3 Corpus Size
As we built TUBELEX using a fixed size (120,000) sample of videos for each language, the final corpus size depends on the number of valid videos after cleaning (see Table 1). The Chinese (18M tokens) and Indonesian (38M tokens) corpora are substantially smaller than the others (163M to 171M tokens). We therefore expect that improvements in correlation could be made by collecting larger corpora, particularly for these two languages, although the effect could be difficult to assess on available data. Increased corpus size also likely benefit language models (see Appendix C).
Sizes of all corpora and datasets used in this study can be found in Appendix E and in Appendix G, respectively.
@@ 7 Conclusion
The data that could be released is limited by copyright, as detailed in the following section.
@@
We only downloaded subtitles for videos that could be found using the YouTube website search function. The search function is restricted to public videos, excluding any unlisted or private videos.121212 https://support.google.com/youtube/answer/157177?hl=en
@@
As we have accessed YouTube without sign-in, our corpus does not contain any subtitles for age-restricted videos, which YouTube defines as not appropriate for viewers under 18.131313 https://support.google.com/youtube/answer/2802167?hl=en Note that while YouTube’s age restriction also applies to “excessive profanity”, some subtitles in our corpus still contain vulgar or otherwise inappropriate language, which we did not attempt to remove.
@@ Appendix C Evaluation of Word Embeddings
The TUBELEX embeddings were trained using the same hyperparameters (see Appendix A) as the Wikipedia embeddings by Grave et al. (2018), while van Paridon and Thompson (2021) used a different setup. All compared embeddings are fastText, but we do not use character -grams to embed out-of-vocabulary words. For fairness, we always evaluate the whole dataset including for out-of-vocabulary words. We only evaluate on English and Spanish, for which comparable evaluation data and pre-trained OpenSubtitles2018 embeddings are available.
@@
Appendix D Details of the Evaluated Corpora

### v2

@@ 1 Introduction
Word frequency is crucial for psycholinguistic research, as well as for assistive or educational applications involving production or comprehension of words.
Psycholinguistic analyses of the relative strength of variables affecting lexical processing, such as word frequency and age of acquisition (e.g. Garlock et al., 2001), hinge on accurate data for these variables. As word frequency and age of acquisition are correlated with each other, having less representative frequency data can easily change the result of such an analysis.
Traditionally, written corpora have been used for estimates of word frequency, with Kučera and Francis (1967) frequency norms long dominating psycholinguistic research of English. While the size of written language corpora has grown over time, speech corpora are still costly to develop and comparably limited in extent. When user-generated text became available on a large scale, it was possible to approximate everyday language exposure by collecting English text from Usenet newsgroups (Burgess and Livesay, 1998), and later French film and TV subtitles (New et al., 2007). Subtitle-based norms for US English, SUBTLEX-US, (Brysbaert and New, 2009) were found more predictive of lexical decision times (LDT) than frequencies based on traditional written corpora or the Usenet-based corpus.
@@
As full corpus data cannot be published due to copyright, we release two basic language models based on the TUBELEX corpus for each language in addition to the frequency lists: a statistical language model (Heafield et al., 2013), which provides smoothed frequencies of word 1-grams to 5-grams, and fastText word embeddings (Bojanowski et al., 2017) to enable modeling of semantic similarity or analogy, as well as representation of words in downstream application. FastText extends the Word2vec model (Mikolov et al., 2013). Preprocessing details and hyperparameters are provided in Appendix A, model sizes in Appendix B, and evaluation of the embeddings in Appendix C.
@@ 2.1 Subtitle Corpora
All of these corpora aim to approximate spoken language and most of them were evaluated against psycholinguistic data. Other subtitle corpora were built for different purposes:
@@ 3.2 Cleaning and Duplicate Removal
We identify the language of each subtitle line using the compressed fastText language identification model444https://fasttext.cc/docs/en/language-identification.html (Joulin et al., 2016a, b). We discard files containing less than 95% of the target language. From the remaining files, we remove both lines that do not contain any valid characters for the target language (e.g. Latin alphabet for English), and lines that are identified as a different language. Lastly, we discard any files less than three lines long. The resulting numbers of files (i.e. files not discarded during cleaning out of the sample of size 120,000) are listed as Cleaned in Table 1.
@@ 3.4 Masking Personal Information
We also use special tokens to replace sequences of digits (after tokenization) and anonymize email addresses, web addresses including those without an explicit protocol (e.g. x.com/username), and apparent social network handles starting with @. Our approach to anonymization is informed by the analysis by Subramani et al. (2023) and extends the approach of Soldaini et al. (2024) by masking web addresses and social network handles.
@@ 3.5 Tokenization and Frequency Lists
English, Indonesian, and Spanish segmented using Stanza (Qi et al., 2020) tokenize,mwt pipeline; Japanese segmented using MeCab (Kudo et al., 2004) with UniDic 2.1.2 (Den et al., 2007, distributed as unidic-lite777https://pypi.org/project/unidic-lite/); Chinese segmented using the jieba888https://github.com/fxsjy/jieba segmenter 0.42.1.
@@
| Status | Evaluation Result | Chinese | English | Indonesian | Japanese | Spanish |
| OK | Subtitles match speech in the target language | 65% | 91% | 84% | 84% | 89% |
| Problematic | Subtitles match song in the target language | — | 2% | 1% | 1% | 2% |
| Subtitles provide audio description (e.g. phone rings) | — | 1% | 1% | 0% | 0% | |
| Audio is neither speech or song | — | 2% | 0% | 0% | 2% | |
| Audio is synthesized speech | 34% | 3% | 13% | 14% | 5% | |
| Audio language differs from the target language | 1% | 1% | 1% | 1% | 2% | |
| Subtitle language differs from the target language | — | — | — | — | — |
@@ 4 Human Evaluation
We performed human evaluation to verify how representative the corpus is of the target languages, and spoken language in particular. We took a sample of size 300 for each language, consisting of 100 videos with 3 random subtitle time stamps for each. The videos were selected using stratified sampling by category and duration999 We divided the videos in three similarly large duration classes: [0, 3 min), [3 min, 10 min), and [10 min, ). for each language. Each sampled timestamp was examined and labeled by a CEFR C2-level non-native speaker for English and by native speakers for the other languages. We examined each example by playing the video and comparing the subtitle cue (the subtitle text displayed at the given timestamp) with the corresponding audio, extending the examined video segment as deemed necessary by the evaluator to categorize the example.
The evaluation results are shown in Table 2. To be evaluated as “OK”, the subtitles must match human speech in audio and both must be in the target language. Otherwise, the instance is assigned to one of the “Problematic” evaluation results. While in principle problems not listed in the table may occur (e.g. the language matches, but the content of audio and subtitles differs) or the listed problems could co-occur, it did not happen in our sample.
Most importantly, among the 1,500 subtitle cues, we have not found a single one whose language would differ from the target language. We have, however, observed different dialects or varieties of each language, as well as apparent non-native speech, sometimes co-occurring in the same video. In the case of Chinese, 2 of the 100 videos were in Cantonese, with traditional Chinese subtitles, while the majority was in Mandarin Chinese. To better understand the composition of the Chinese subtitles, we also analyzed the script used in the whole corpus, and found that 66% videos use simplified Chinese, 33% videos use traditional Chinese, and 1% mix both.101010 We used the Hanzi Identifier package (https://github.com/tsroten/hanzidentifier). and considered subtitles mixed if they contained the non-majority script variant on at least 2 lines and at least 5% of lines.
@@ 5 Extrinsic Evaluation
Note that the datasets available for different languages generally have different characteristics (such as part of speech or word frequency distribution), so while our experiments allow comparison of different corpora for a particular task and language, they do not allow comparison across languages. For instance, we should not surmise that TUBELEX provides better data for Spanish than for English (or that lexical decision time is generally more strongly correlated with frequency in Spanish than in English) only because TUBELEX achieves (or because all corpora achieve) a higher correlation for the task in Spanish than English.
@@ 5.1 Evaluated Corpora and Resources
GINI, a Twitter-based metric of words’ dispersion in frequency of use by different people (Murayama et al., 2018; also see Appendix D); Wikipedia; wordfreq, a Python library (Speer, 2022) pooling frequency from multiple corpora. Wordfreq combines Wikipedia, Twitter and a subtitle corpus for each of the evaluated languages, as well as 4 more sources for English, Chinese, and Spanish, and 2 more for Japanese. The subtitle data used by wordfreq is OpenSubtitles2018, SUBTLEX-US and SUBTLEX-UK for English, SUBTLEX-CH for Chinese.
For each corpus, we provide technical details in Appendix E, and token and type counts in Appendix F.
@@ 5.2 Computing Frequency
where is the number of occurrences of the word , #tokens is the total number of tokens in the corpus, and #types is the number of types in the corpus. As a result, even words missing in the corpus are assigned a non-zero frequency. In all experiments we use the logarithm of frequency. Appendix E provides details about specific corpora.
@@ 5.3 Lexical Decision Time
Other corpora either performed comparably, but only covered a single language (SubIMDB, EsPal), or underperformed noticeably in at least one language (SUBTLEX in Spanish, wordfreq in Chinese, Wikipedia in all languages, and GINI and ACTIV-ES in the one language they cover). Furthermore, Figure 1 shows that TUBELEX and SUBTLEX perform remarkably well relative to their size.
@@ 5.4 Word Familiarity
We measure correlation (PCC) with mean word familiarity from five databases: Chinese familiarity ratings (Su et al., 2023), English MRC lexical database (Coltheart, 1981; Coltheart and Wilson, 1987), Indonesian lexical norms (Sianipar et al., 2016), Japanese word familiarity ratings (Asahara, 2019)111111 We use the published ratings for reception estimated using a Bayesian linear mixed model. , and Spanish lexical norms (Guasch et al., 2016). Evaluation on three alternative, smaller databases for English, Spanish, and Japanese can be found in Appendix G.
As shown in Table 4, in Japanese, TUBELEX’s correlation is the strongest one, and in all other languages TUBELEX’s correlation is either the strongest one or not significantly weaker. Correlations without any significant difference from TUBELEX are achieved by SUBTLEX in Chinese, by all subtitle corpora in English, by wordfreq in Indonesian, and by all subtitle corpora except EsPal, and by CREA-Spoken in Spanish. In Japanese, GINI achieves a remarkably strong correlation, significantly stronger than TUBELEX but still significantly weaker than TUBELEX (∗∗) and TUBELEX (∗∗∗).121212We computed the levels of significance separately, as Table 4 compares all corpora only against TUBELEX.
@@ 5.5 Lexical Complexity
As the dataset was used for evaluation of lexical complexity prediction in a shared task (Shardlow et al., 2024), we also compare predictions based on TUBELEX with top shared task participants. To do so, we fit a linear regression model using a single variable, log-frequency or GINI values to the shared task’s trial data (30 instances for each language), and clip the predicted values to the range [0, 1]. We compare our results with the shared task submissions that achieved the highest coefficient of determination (and also the highest correlation) in individual languages (TMU-HIT, Enomoto et al., 2024; and GMU, Goswami et al., 2024), and on average across all shared task languages (Archaeology, Cristea and Nisioi, 2024).
As shown in Table 6, the best results are achieved by TUBELEX, closely followed by TUBELEX, in English; TUBELEX, closely followed by TMU-HIT and TUBELEX, in Japanese; and by TMU-HIT in Spanish, where it outperformed others by a large margin. As we are using cited values achieved by task participants, we cannot evaluate statistical significance in this case.
@@
It should be noted that, in this case, we are evaluating the prediction of lexical complexity using , not a mere correlation with lexical complexity.131313 While can also be defined as the square of PCC, we use the definition consistent with the shared task evaluation (Shardlow et al., 2024), and implemented in scikit-learn (https://scikit-learn.org/) as r2_score. With this definition, is a linear function of the mean squared error, therefore penalizing misprediction of mean and variance.
If we looked only at correlation, thus ignoring misprediction of mean and variance, TMU-HIT’s predictions would be more strongly correlated than TUBELEX’s in the three languages, and those of the other two systems in English (complete results provided in Appendix I). This may indicate limitations of LLM prompting as a regression method.
@@ 6.1 Tokenization and Lemmatization
For Japanese, the base form performs the best for familiarity, and default tokenization for non-native lexical complexity. Both always outperform the orthographically normalized lemma, which show the importance of the exact written form in Japanese.
@@ 6.3 Corpus Size
As we built TUBELEX using a fixed size (120,000) sample of videos for each language, the final corpus size depends on the number of valid videos after cleaning (see Table 1). The Chinese (18M tokens) and Indonesian (38M tokens) corpora are substantially smaller than the others (163M to 171M tokens). We therefore expect that improvements in correlation could be made by collecting larger corpora, particularly for these two languages, although the effect could be difficult to assess on available data. Increased corpus size would also likely benefit language models (see Appendix C).
Sizes of all corpora and datasets used in this study can be found in Appendix F and in Appendix H, respectively.
6.4 Beyond Frequency
Our goal was to evaluate the TUBELEX corpus as an approximation of spoken vocabulary. For a comprehensive comparison with other corpora, we limited the evaluation to word frequency – the only statistic available for many of the compared corpora. Similarly, most of the data we used for evaluation (LDT, familiarity and complexity datasets) is limited to or focused on single-word items. There are several ways TUBELEX could be used or extended for other purposes:
Language modeling.
More complex language models would allow further applications of the corpus data such as modeling human surprisal (e.g. Wilcox et al., 2023). We used the current corpus data to train two basic language models: an -gram model and word embeddings, which achieved a mixed performance in our evaluation (Appendix C). Collecting larger data may improve the performance of the embeddings and it would be essential for training more complex language models (masked language models or generative language models). We expect that, similarly to the results achieved by the current embeddings, models trained exclusively on subtitles would lack in some areas, and consequently that training on mixed data would be suitable for a wide range of applications.
Measures of dispersion and contextual diversity.
Single-word corpus statistics are not limited to frequency. We have compared TUBELEX frequency to GINI, a measure of dispersion based on Twitter data, but we leave evaluation of various dispersion metrics computed from TUBELEX for future research. Computation of dispersion metrics generally requires a corpus divided into suitable units such as documents, which TUBELEX provides (videos and channels). Another metric that uses such units is contextual diversity (the number of units in which a word occurs), proposed by Adelman et al. (2006) as an alternative to frequency for psycholinguistic modeling. TUBELEX word lists readily provide contextual diversity as numbers of videos and channels for each word.
Combining sources.
Combination of frequencies from multiple corpora often achieve more robust results than using a single corpus. While TUBELEX often outperformed wordfreq, which is a resource combining multiple corpus frequencies, combining TUBELEX frequencies with other sources (e.g. film subtitles) could also result in a more robust performance, especially for languages underrepresented on YouTube.
@@ 7 Conclusion
TUBELEX data can be easily used in applications that require spoken vocabulary frequencies (examples given in Section 2.2), which have typically relied on film subtitles but lacked a suitable resource for many languages. Our method can be used to create TUBELEX corpora for additional languages as well. We see extension beyond using single-word frequencies (outlined in Section 6.4) as a promising direction for future research.
@@
Estimating how much subtitle data for a particular language is available on YouTube requires non-trivial effort. For a reliable estimate, it is necessary to identify videos by keywords search, scraping the metadata for each video, and evaluating at least a sample of the subtitles using automated language identification (see Section 3). We did this for two additional languages: For Czech, which has much fewer speakers than any of the current five languages (9.6 million L1 speakers; Eberhard et al., 2024), we found enough data to build a corpus comparable to the current five TUBELEX language. We could not, however, find enough data for Pashto with 44 million L1 speakers (ibid.). We hypothesize that this reflects not only the number of speakers and the popularity of YouTube among them, but also the relative prestige of Pashto in the two multi-lingual countries where it is mainly spoken, Afghanistan and Pakistan. Similar challenges may also affect other low-resource languages.
@@
We only downloaded subtitles for videos that could be found using the YouTube website search function. The search function is restricted to public videos, excluding any unlisted or private videos.141414 https://support.google.com/youtube/answer/157177?hl=en
@@
As we have accessed YouTube without sign-in, our corpus does not contain any subtitles for age-restricted videos, which YouTube defines as not appropriate for viewers under 18.151515 https://support.google.com/youtube/answer/2802167?hl=en Note that while YouTube’s age restriction also applies to “excessive profanity”, some subtitles in our corpus still contain vulgar or otherwise inappropriate language, which we did not attempt to remove.
@@ Appendix C Evaluation of Word Embeddings
The TUBELEX embeddings were trained using the same hyperparameters (see Appendix A) as the Wikipedia embeddings by Grave et al. (2018), while van Paridon and Thompson (2021) used a different setup. All compared embeddings are fastText, but we do not use character -grams to embed out-of-vocabulary words. For fairness, we always evaluate the whole dataset including out-of-vocabulary words. We only evaluate on English and Spanish, for which comparable evaluation data and pre-trained OpenSubtitles2018 embeddings are available.
@@
Appendix D GINI Metric
The GINI metric was proposed by Murayama et al. (2018) for simplification and readability assessment. It is inspired by the Gini index of income inequality, and measures words’ dispersion in frequency of use by different people. Similarly to TUBELEX, it uses user-generated data (from Twitter), and its pre-computed values are available for English and Japanese. As the original study is available only in Japanese, we summarize its computation for the reader’s convenience:
1.
Construct a matrix of numbers of word occurrences shaped (#users, #words).
2.
Normalize twice: first making each row sum to 1, then making each column sum to 1.
3.
For each column (word) , compute:
4.
To compute the final values of the metric, apply: .
Appendix E Details of the Evaluated Corpora
Table 11 shows details of the evaluated corpora. We compute frequency with Laplace smoothing (see Section 5.2) from them with the following exception: for ACTIV-ES and wordfreq, which do not provide token counts, we instead assign the corpus minimum frequency to missing words. We also directly use GINI values, analogously assigning the corpus maximum value to missing words, as high GINI values indicate high dispersion. The words featured in our experiments may be tokenized as multiple tokens. We assign them the minimum of the frequencies of the individual tokens, and maximum of the GINI values. In all experiments, we use the logarithm of frequency. For GINI, we use the additive inverse () of the original metric’s values described in Appendix D, for the purpose of comparison with log-frequencies used for the other corpora.